OpenStudy (anonymous):

Can someone show me how to do this? Solve the inequality and graph the solution set. -4a-8b< or = 6a+22

OpenStudy (anonymous):

\[-4a-8b\leq 6a+22\] \[-8b\leq 6a+4a+22\] \[-8b\leq 10a+22\] \[-8b-22\leq 10a\] \[\frac{-8b-22}{10}\leq a\] \[\frac{-5b-11}{5}\leq a\] \[-\frac{b+11}{5}\leq a\]

OpenStudy (anonymous):

have I told you guys lately how much I love you and appreciate you????

OpenStudy (anonymous):

\[\color{#ff0099}{\huge{\text{smile :)}}}\]

OpenStudy (anonymous):

do u always divide by the number with the variable?